Solutions to Radical Equations

Explain why possible solutions to a radical equation must be checked to determine the actual solutions.
Answer #1

Possible solutions have to be checked because when we are solving radical equation we are squaring the terms to remove radical sign, due to this negative numbers will become positive and solution will not satisfy the original equation.

Hence, possible solutions should be checked by putting in original equation to get actual solution

For example √(a-5) = -2

Squaring both sides

a-5 = 4 or a = 9

But when we put a= 9 in original equation result will be

√(9-5) = 2

Which is not equal to -2 . Hence, possible solution need to be checked.

If you like my answer please give me good rating thank you